### v2

# SDE-SQL: Enhancing Text-to-SQL Generation in Large Language Models via Self-Driven Exploration with SQL Probes

Abstract
Recent advances in large language models (LLMs) have led to substantial progress on the Text-to-SQL task. However, existing approaches typically depend on static, pre-processed database information supplied at inference time, which restricts the model’s capacity to deeply comprehend the underlying database content. In the absence of dynamic interaction, LLMs are limited to fixed, human-curated context and lack the ability to autonomously query or explore the data. To overcome this limitation, we introduce SDE-SQL, a novel framework that empowers LLMs to perform Self-Driven Exploration of databases during inference. This is achieved through the generation and execution of SQL probes, enabling the model to actively retrieve information and iteratively refine its understanding of the database. Unlike prior methods, SDE-SQL operates in a zero-shot setting, requiring no in-context demonstrations or question-SQL pairs. Evaluated on the BIRD benchmark with Qwen2.5-72B-Instruct, SDE-SQL achieves an 8.02% relative improvement in execution accuracy over the vanilla Qwen2.5-72B-Instruct baseline, establishing a new state-of-the-art among open-source methods without supervised fine-tuning (SFT) or model ensembling. Furthermore, when combined with SFT, SDE-SQL delivers an additional 0.52% performance gain.
SDE-SQL: Enhancing Text-to-SQL Generation in Large Language Models via Self-Driven Exploration with SQL Probes
Wenxuan Xie1, Yaxun Dai2, Wenhao Jiang3thanks: Corresponding author 1South China University of Technology, 2Soochow University 3Guangdong Laboratory of AI and Digital Economy (SZ) lancelotxie601@gmail.com, cswhjiang@gmail.com
1 Introduction
Text-to-SQL is a long-standing task in natural language processing that focuses on translating natural language questions into executable SQL queries. This capability not only empowers non-expert users to interact with structured databases seamlessly, but also mitigates hallucination issues in question-answering systems by grounding responses in factual, database-stored information.
Recent advances in large language models (LLMs) have led to significant improvements in the performance and accuracy of Text-to-SQL systems. LLM-based approaches have surpassed 90% execution accuracy on the original Spider dataset (Yu et al., 2019), and have demonstrated promising results on more complex and diverse benchmarks such as BIRDLi et al. (2023). Despite these advances, a noticeable gap remains between current model performance and human-level capabilities—particularly on the recently introduced Spider 2.0 benchmark (Lei et al., 2025), which poses more realistic and challenging scenarios for semantic parsing.
Figure 1: The Workflow of SDE-SQL, which consists of three parts: 1) Schema Linking: which retrieves and selects useful database schema; 2)SQL Generation: performing zero-shot SQL generation based on two-phase self-driven exploration; 3)SQL Refinement: which refines the SQL with the execution results of the Sub-SQLs and SQL Probes.
Contemporary large language model (LLM)-based approaches to Text-to-SQL typically comprise three core components: schema linking, SQL generation, and SQL refinement. In the schema linking stage, prior work has primarily focused on aligning natural language questions with relevant database schema elements, improving precision and contextual relevance. During SQL generation, various methods have been proposed to decompose complex questions and incorporate reasoning strategies. In the refinement stage, the categorization of SQL error types has become more systematic, enabling the development of targeted correction mechanisms.
Despite these advances, one crucial aspect of SQL remains largely underexplored: its inherent interactivity as a database interface that supports fast and informative execution. This underutilized property may partially account for the performance gap between LLM-based systems and human experts.
To address this, we propose SDE-SQL, a novel framework that incorporates Self-Driven Exploration into both the generation and refinement stages, as illustrated in Figure 1. In addition to generating the final SQL query that directly answers the natural language question, the model autonomously generates and executes a sequence of auxiliary queries—termed SQL Probes—designed specifically to explore and extract informative signals from the database.
For schema linking, we leverage entity-based techniques including value retrieval and soft linking. During the generation phase, the model engages in a two-stage exploration process based on the question and schema, enabling it to iteratively refine its understanding of the database content and perform zero-shot reasoning grounded in the retrieved information.
Following generation, we incorporate a two-stage exploration process into the refinement phase. For SQL queries that return explicit execution errors, the model directly revises them based on the error feedback. For queries that execute successfully but return empty results, the first stage of exploration uses the execution results of decomposed sub-queries (Sub-SQLs) to help the model diagnose the underlying issue. In the second stage, the model generates targeted SQL Probes to explore possible solutions, and selects the most promising one to produce the final refined query.
Empirically, SDE-SQL achieves an execution accuracy of 67.67% on the BIRD benchmark using Qwen2.5-72B-Instruct in a zero-shot setting. With supervised fine-tuning (SFT), the performance further improves to 68.19%.
Our main contributions are as follows:
  • We propose SDE-SQL, a novel framework that leverages Self-Driven Exploration to enhance the reasoning and interaction capabilities of LLMs in the Text-to-SQL task, significantly narrowing the gap with human experts.
  • We introduce a unified exploration mechanism across both SQL generation and refinement stages, enabling LLMs to actively query the database, diagnose potential errors, and iteratively improve query quality.
  • We conduct extensive experiments on the BIRD and Spider benchmarks, along with ablation studies, validating the effectiveness of Self-Driven Exploration.
  • We build a small-scale dataset for supervised fine-tuning (SFT) on exploration and generation tasks, and show that targeted module-level fine-tuning further improves the performance of SDE-SQL.
2 Related Work
Transforming natural language questions into database queries is a classic task, the earliest works used inductive logic programming and human-designed templates to accomplish this task(Zelle and Mooney, 1996). In recent years, the advancement of Text-to-SQL technologies can be broadly categorized into two stages, driven by progress in natural language processing.
2.1 Traditional Seq2Seq Model-Based Methods
Previous work primarily focused on improving encoding or decoding methods, as the seq2seq model framework consists of two main components, the encoder and the decoder. IRNet employed a bidirectional LSTM to encode the question and a self-attention mechanism to encode the database schema, ultimately using an LSTM as a grammar-based decoder(Guo et al., 2019). In order to effectively capture the relationship between the database schema and the question, RAT-SQL develops an encoder with a relation-aware self-attention mechanism(Wang et al., 2020). After that, Cai et al. (2022) and Cao et al. (2021) utilized graph neural networks to encode the relationships between the schema and the query. Leveraging the exceptional capabilities of pre-trained language models (PLMs) across various NLP tasks, Hwang et al. (2019) was the first to incorporate BERT as its encoder. For improvements in the decoder, Xu et al. (2017) and Choi et al. (2020) focused on sketch- based decoding method. To reduce time consumption during inference, SDSQL presented the Schema Dependency Learning and removed execution-guided (EG) decoding strategy(Hui et al., 2021).
2.2 LLM-Based Methods
With the advent of LLMs, the Text-to-SQL field has experienced a groundbreaking innovation, bringing about significant changes in the approach to the task.
Methods Based on Prompt Engineering
Rajkumar et al. (2022) evaluated the potential of LLMs in the Text-to-SQL task, demonstrating the remarkable capability of LLMs in this task. Building on in-context learning, DAIL-SQL (Gao et al., 2023) introduced a novel prompt engineering approach that improves the Text-to-SQL performance of LLMs through question representation, demonstration selection, and demonstration organization. Based on Chain-of-Thought(CoT) reasoning style(Wei et al., 2023), DIN-SQL(Pourreza and Rafiei, 2023), Divide-and-Prompt(Liu and Tan, 2023), CoE-SQL(Zhang et al., 2024a) and SQLfuse(Zhang et al., 2024b) designed CoT templates with reasoning steps in the prompt to elicit chain thinking. To enhance the ability of LLMs in handling complex problems, QDecomp(Tai et al., 2023), DIN-SQL(Pourreza and Rafiei, 2023), MAC-SQL(Wang et al., 2025) and MAG-SQL(Xie et al., 2024) decomposed complex natural language questions and solve them step by step. Besides, MCS-SQL(Lee et al., 2024), CHASE-SQL(Pourreza et al., 2024a) and CHESS(Talaei et al., 2024) enhanced performance by generating a large set of candidate SQL queries during the inference stage and selecting the most suitable ones.
Methods Based on Fine-tuning
Although prompt engineering methods based on closed-source models, like GPT-4o(OpenAI et al., 2024), perform well in the Text-to-SQL task, they face issues such as high costs, inability to guarantee privacy, and limited flexibility. Therefore, fine-tuning open-source models for the Text-to-SQL task holds significant practical value and application potential. DTS-SQL(Pourreza and Rafiei, 2024) and SQLfuse(Zhang et al., 2024b) explored fine-tuning LLMs for both schema linking and SQL generation. SQL-PaLM(Sun et al., 2024), Open-SQL(Chen et al., 2024), XiYan-SQL(Gao et al., 2025) and CodeS(Li et al., 2024) fine-tuned open-source LLMs on carefully selected data, while CodeS specifically adopted an incremental pre-training approach using a specially curated SQL-centric corpus. In addition, there are some novel perspectives. DELLMHong et al. (2024) specifically fine-tuned a Data Expert Language Model that provides domain knowledge, while SQL-GENPourreza et al. (2024b) proposed a novel Mixture-of-Experts (MoE) architecture to handle multiple SQL dialects.
3 Methodology
3.1 Entity-based Schema Linking
In the Text-to-SQL task, schema linking refers to the process of identifying and selecting the relevant tables, columns, and values from the database based on the input natural language question. To improve the accuracy of linking, we use an entity-based linking approach, including Value Retrieval and Soft Schema Linking.
3.1.1 Entity-based Value Retrieval
Similar to the retrieval module in Talaei et al. (2024), we first employ an LLM to extract entities from the natural language question through few-shot learning. And then the value retriever identifies similar values in the database based on Locality Sensitive Hashing (LSH) and semantic similarity.
3.1.2 Entity-based Soft Schema Linking
To improve the tolerance in the schema linking stage, we chose the soft schema linking method, like the approach in Xie et al. (2024). We employ a one-shot manner to prompt LLM to select the relevant columns based on each entity. For the selected columns, we provide as much detailed information as possible during the subsequent SQL generation, including the column name, type, column description, value examples, and value descriptions. For the unselected columns, we only retain the column name and type. This approach not only significantly reduces the input length, allowing the language model to focus on the most relevant database schema during generation, but also enhances tolerance by preventing the removal of useful columns that were not chosen.
3.2 Generation Based on Self-Driven Exploration
In previous Text-to-SQL research, SQL has often been viewed primarily as an intermediate result or final output, with its inherent functionality mostly overlooked. Therefore, we introduce the concept of SQL Probes. SQL Probes, literally meaning SQL queries that function as probes, are specifically designed for exploring the database based on current natural language question. Formally, we define the task as a mapping from a natural language query $Q$ and a database schema $D$ to a corresponding SQL query $S$. The natural language query $Q$ is composed of two parts: the target and the conditions (Xie et al., 2024). Typically, the target corresponds to the main SELECT clause in the SQL query $S$, while the conditions correspond to the other clauses in $S$, such as the WHERE clause. Figure 2 is an example.
Figure 2: An example of Text-to-SQL.
To obtain a specific SQL representation, entities must first be mapped to the corresponding columns and values in the database. Whether this step can be executed accurately depends on how well the language model understands the database.
However, the information provided by the previously processed database schema is far from sufficient. Real-world databases are often highly complex and messy. Different tables may contain many columns with the same meaning (representing the same item), and the values in these columns might have different formats, with some values even existing only in specific tables. In the absence of sufficient information, LLM can only randomly identify combinations from these similar columns and values. This is also one of the key reasons behind the LLM’s especially unstable performance in this task. During evaluation, it is frequently observed that the model can correctly predict some examples at times, while failing on the same examples at other times. In prior work, some methods have involved generating multiple SQL-candidates with the language model, followed by selecting the most appropriate one. Nevertheless, this approach fails to address the underlying problem. We propose that the most fundamental solution is to empower LLM with the ability to dynamically interact with the database. In SDE-SQL, LLM performs a two-stage self-driven exploration within the database before generation.
Figure 3: Condition SQL Probes Generation Process Illustrated Using a Tree Structure.
3.2.1 Candidates Exploration
The goal of this stage of exploration is to enable the large language model to query the database for information regarding both the Targets and a single Condition, and then select appropriate candidates for each target and condition. Since an entity in a natural language question is mapped to either a column or a value (or both a column and a value) in the database, LLM needs to determine the candidate columns and candidate values for each entity. Initially, the language model generates several Base SQL Probes, which enumerate candidate columns for the Targets. These SQLs focus solely on querying the Targets without any additional conditions. Following this, Condition SQL Probes are created, where each Probe typically extends a Base SQL Probe by adding a column candidate and maybe a value candidate corresponding to a specific condition. Assuming each set of candidates contains two options, the generation of Condition SQL Probes is illustrated in Figure 3, where each root-to-leaf path corresponds to a specific Condition SQL Probe. We refer to the condition description of each Condition SQL Probe as a Condition Description Candidate. For example, in Figure 3, one Condition Description Candidate is:
SELECT Phone FROM users WHERE name = ’JohnCharlieHinton’ AND location = ’Uni%tedKingdom’;
3.2.2 Combinations Exploration
Based on the results of the previous stage’s exploration, the scope of candidates has been narrowed down. Now, it is necessary to combine all the conditions to find the most suitable candidate combination. For SQL queries that return no results, the corresponding candidate combination is definitely unsuitable.
3.2.3 Zero-shot Generation with Exploration Results
In our experiments, we found that existing methods do not fully leverage the large language model’s potential for SQL generation. For example, strategies such as designing new decomposition approaches to allow the model to progressively solve complex problems, using various prompt techniques to generate multiple candidates for selection, or employing search strategies like Monte Carlo tree search(MCTS) to enhance the inference capability of language models, can lead to modest improvements in model performance. However, these gains are still significantly smaller than those achieved by providing the model with sufficient information.
Therefore, in SDE-SQL, the LLM generator generates SQL based on the database schema and the results from the previous two exploration stages, without relying on any question-SQL pairs as few-shot examples or using any question decomposition strategies. To improve the accuracy and robustness of SQL generation, we adopt a self-consistency strategy that selects the most consistent answer by comparing the execution results of multiple generated SQL queries.
Figure 4: An illustration of the proposed refinement process with exploration in SDE-SQL.
3.3 Refinement Based on Self-Driven Exploration
In the past, existing techniques based on In-Context Learning have introduced detection and repair solutions for Text-to-SQL errors, with each solution differing in its approach to error identification algorithms and the supplementary data provided to assist LLM in comprehending and rectifying these errors.(Shen et al., 2025)
For Syntax errors and Schema errors, the error feedback after execution already contains sufficient information, allowing LLMs to effectively complete the correction of SQL. However, for some other more complex errors, they typically result in empty query results without any error messages. Even when humans attempt to correct these errors, they cannot do so in one go; instead, they need to write some SQL statements for debugging and diagnose the problem based on the execution results of these queries. The current approach involves continuously regenerating until the repairs is successful or the attempt limit is reached. Throughout this repair process, LLM does not receive any useful information, and its reasoning abilities are not fully utilized. In other words, the reason for the error is never identified.
Therefore, in SDE-SQL, we introduce a comprehensive Self-Driven Exploration phase prior to SQL revision. For queries that yield empty results, the refinement process is divided into three distinct stages: the Error Cause Identification Stage, the Solution Exploration Stage, and the Modification Stage, as illustrated in Figure 4.
3.3.1 Error Cause Identification Stage
In a complex SQL statement, multiple tables may be involved and multiple conditions may be applied simultaneously, making it difficult to pinpoint the issue by directly analyzing the entire SQL. Therefore, we need to conduct a fine-grained diagnosis. To generate a series of Sub-SQLs as SQL Probes for diagnostic purposes, we developed a decomposer based on SQLGlot. The decomposer first converts complex SQL queries into Abstract Syntax Trees (ASTs) and then identifies indivisible condition units by analyzing node types and their relationships. These identified subtrees within the AST serve as the foundation for generating semantically valid Sub-SQLs, and the execution results of these Sub-SQLs will be provided to LLM to assist it in accurately diagnosing and pinpointing issues in the original query. An example of the decomposition result is shown in Figure 8.
3.3.2 Solution Exploration Stage
We have summarized five possible reasons that may lead to an empty query result. At this stage, LLM need to analyze the probe results in the previous stage to derive hypotheses about possible error causes, and then generate a series of SQL probes to assist in exploring potential solutions to these errors.
Conditions conflict or condition duplication
This error refers to situations where data can be found when executed under a single condition, but when multiple conditions are combined, no data that meets the requirements can be found (resulting in an empty query result). There are two possible reasons for this error: conflicting combinations of multiple different conditions or redundant descriptions of a single condition using different columns. (i) Conditions conflict typically arises when an entity in a condition corresponds to multiple possible candidate columns, and only a specific candidate column combined with other conditions can yield the corresponding data item. An example is shown in Figure 5. (ii) Condition duplication occurs when an entity in a condition maps to multiple candidate columns, causing the SQL generated by the large language model to inadvertently employ these various candidate columns in describing the same condition, ultimately resulting in the failure to retrieve data that fulfills the intended condition.
Unnecessary Table Joins
The SQL may include unnecessary table joins, resulting in no records satisfying the conditions in the final intersection.
Mismatch between column and value
This error arises when either the value format does not match the selected (but correct) column, or when a similar-looking column is chosen that does not contain the intended value.
Sub-query Scope Inconsistency
Sometimes, the scope of the sub-query may be inconsistent with that of the main query, especially when using the MIN/MAX functions in the sub-query, which often leads to an empty query result. Figure 6 shows an example that the row of data retrieved in the sub-query does not exist in the result after the JOIN of these two tables.
3.3.3 Target Checking After Refinement
For an SQL query, the most important part is actually the target of the query, which refers to the columns being selected in the SELECT clause. If the query target in SQL does not align with the original query target in the natural language question, then the transformation is undoubtedly a failure. However, when LLMs generate SQL, they sometimes include columns that are not part of the query target in the SELECT clause, such as columns used in the conditions. Therefore, after refining the SQL based on the execution results, it is necessary to check whether the query target in the SQL matches the query target in the natural language question. To avoid introducing new errors at this stage, we allow the large language model to only determine if unnecessary target columns are selected in the SQL. If such columns are found, they will be removed without affecting the execution. The procedure is illustrated in Figure 7.
3.4 Supervised Fine-Tuning (SFT)
To further enhance the model’s ability to autonomously explore the database and utilize exploration results to generate more accurate SQL, we also perform supervised fine-tuning (SFT) on the model. The training data is sampled from the training set of BIRD. We employed a prompt-based pipeline to roll out data, and the examples that eventually produced correct SQL were regarded as valid data for fine-tuning the model.
Among the 9,428 data points, 5,231 valid samples were obtained through sampling. From the reasoning trajectory of each example, we extract two components: (i) the exploration phase, where SQL probes are generated; and (ii) the prediction phase, where the final SQL query is generated based on the exploration results.
4 Experiments
In this section, we first introduce the experimental setup, and then report and analyze the results.
4.1 Experimental Setup
4.1.1 Dataset and Metrics
Spider Yu et al. (2019) is a widely adopted benchmark dataset for the Text-to-SQL task. It is large-scale, cross-domain, and complex, containing 10,181 natural language questions and 5,693 corresponding SQL queries across 200 different databases. As a challenging benchmark of Text-to-SQL task, the recently proposed BIRD dataset Li et al. (2023) includes 95 large-scale real databases with dirty values, featuring 12,751 unique question-SQL pairs. The databases within the BIRD dataset, similar to those in real-world scenarios, exhibit inherent ambiguities. Accordingly, detailed descriptions are provided for each column, along with external knowledge. In this work, we choose Execution Accuracy (EX) as the metric, since it reflects the accuracy of the results returned by the executed SQL queries. This metric considers various SQL formulations for the same question, providing a more precise and fair evaluation of the outcomes.
Method Simple Moderate Challenging All
SDE-SQL + Qwen2.5-72B-Instruct 74.92 57.76 53.10 67.67
   w/o Soft Schema Linker 73.51 58.84 50.34 66.88$\downarrow$0.79
   w/o Exploration Before Generation 72.97 56.46 48.97 65.71$\downarrow$1.96
   w/o Refinement Module 72.97 55.60 48.97 65.45$\downarrow$2.22
    w/o Exploration in Refinement 72.86 56.68 51.72 65.97$\downarrow$1.70
    w/o Target Checking 73.19 57.76 49.66 66.30 $\downarrow$1.37
   w/o Exploration in Generation & Refinement 72.11 54.31 48.28 64.47$\downarrow$3.20
SDE-SQL + Fine-tuned Explorer 74.16 59.26 55.17 67.86$\uparrow$0.19
SDE-SQL + Fine-tuned Generator 74.49 58.19 55.86 67.80$\uparrow$0.13
SDE-SQL + Fine-tuned Explorer & Generator 74.70 58.84 56.55 68.19$\uparrow$0.52
Table 1: Execution accuracy of SDE-SQL on BIRD dev set in the ablation study.
4.1.2 SFT Settings
For both exploration task and generation task, we conducted 24-hour training on 8 NVIDIA A800 GPUs with Qwen2.5-72B-Instruct. The detailed training hyperparameters are provided in Table 4.
4.1.3 Baselines
To enable a comprehensive comparison, we selected representative methods based on closed-source models and representative methods based on open-source models without model ensemble as baselines.
4.2 Main Results
4.2.1 BIRD Results
When evaluated on the BIRD dev dataset, SDE-SQL based on Qwen2.5-72B-Instruct outperforms most GPT-4-based methods and the majority of open-source models, achieving an execution accuracy of 68.19% after fine-tuning, as shown in Table 2. Even in the training-free setting, it achieves a strong performance of 67.67%, further highlighting the effectiveness of our approach.
Method dev(EX)
AskData + GPT-4o 75.36
CHASE-SQL + Gemini 74.46
XiYan-SQL 73.34
OpenSearch-SQL, v2 + GPT-4o 69.30
CHESS 68.31
Distillery + GPT-4o 67.21
MCS-SQL 63.36
MAC-SQL + GPT-4 59.39
DAIL-SQL + GPT-4 54.76
DIN-SQL + GPT-4 50.72
GPT-4 46.35
DTS-SQL + DeepSeek-7B 55.80
SFT CodeS-15B 58.47
SQL-o1 + Llama3-8B 63.4
OneSQL-v0.1-Qwen-32B 64.60
XiYanSQL-QwenCoder-32B 67.01
Qwen2.5-72B-Instruct 60.17
SDE-SQL + Qwen2.5-72B-Instruct 67.67
SDE-SQL (SFT) 68.19
Table 2: The experimental results of competing model on the BIRD dataset.
4.2.2 Spider Results
As shown in Table 3, SDE-SQL fine-tuned solely on the BIRD training set achieves competitive results on Spider benchmark, surpassing GPT-4-based methods and most open-source models, which underscores its strong generalization ability. Nevertheless, the performance gain is relatively modest, as a large portion of SQL queries in the Spider dataset produce empty execution results, thereby limiting the effectiveness of feedback from database exploration.
Method dev(EX) test(EX)
SDE-SQL (SFT) 87.5 88.5
SDE-SQL + Qwen2.5-72B-Instruct 87.3 88.3
MAC-SQL + GPT-4 86.8 82.8
SENSE-13B 84.1 83.5
SQL-o1 + Llama3-8B 87.4 85.4
DAIL-SQL + GPT-4 84.4 86.6
ROUTE + Qwen2.5-14B 87.3 87.1
DIN-SQL + GPT-4 82.8 85.3
GPT-4 (zero-shot) 73.4 -
Qwen2.5-72B-Instruct 73.9 84.0
Table 3: The experimental results of competing model on the Spider dataset.
4.3 Ablation Study
For each module in SDE-SQL, we conduct ablation studies on the development set of BIRD benchmark, which is shown in Table 1. In addition, we evaluate the effect of incorporating the fine-tuned explorer and generator into the pipeline. The results demonstrate that each component plays an important role, with the introduction of the two exploration phases leading to particularly significant performance improvements. Besides, modules fine-tuned on their respective sub-tasks can further enhance the overall performance of the workflow.
5 Conclusion
In this work, we propose SDE-SQL, a novel Text-to-SQL framework that integrates Self-Driven Exploration into both the SQL generation and refinement stages. By enabling LLMs to proactively interact with databases through SQL probes, SDE-SQL bridges the gap between static query generation and dynamic, execution-based reasoning. This exploration mechanism allows LLMs to uncover latent schema semantics and execution patterns, significantly improving their ability to produce executable and semantically accurate SQL queries. Extensive experiments on the BIRD and Spider datasets demonstrate the effectiveness of SDE-SQL, with the model achieving an execution accuracy of 68.19% on BIRD after supervised fine-tuning. Ablation studies confirm the contributions of key components—especially the exploration pipeline and the fine-tuning strategies. As future work, we plan to explore tighter integration of exploration signals into model training to further strengthen the model’s reasoning capabilities.
6 Limitation
Although self-driven exploration significantly enhances the potential of large language models in Text-to-SQL tasks, our current approach has several limitations. In SDE-SQL, database exploration is entirely prompt-driven—meaning that the effectiveness of the exploration process heavily depends on the design and quality of manually crafted prompts. Poorly constructed prompts may lead the model to generate uninformative or redundant SQL probes, thereby limiting its ability to acquire meaningful schema knowledge or execution insights. Moreover, relying solely on prompt engineering can restrict the model’s capacity for deeper reasoning, as it lacks mechanisms for adaptive learning based on feedback from the environment.
Another limitation is the model’s inability to autonomously refine its exploration strategy over time. Since each SQL probe is generated statically from prompts, the model cannot dynamically adjust its behavior based on prior successes or failures during the exploration process. This constraint reduces the overall flexibility and learning efficiency of the system.
To address these issues, future work will focus on making database exploration more intrinsic to the model itself. One promising direction is to incorporate reinforcement learning or other feedback-driven learning paradigms, allowing the model to iteratively refine its probing strategies based on execution outcomes. By enabling the model to learn from its own interactions with the database, we hope to develop a more robust, adaptive framework capable of deeper, context-aware reasoning in complex database environments.
Appendix A Examples of Error Cause
Figure 5: An Example of Condition Conflict
Figure 6: An example of Sub-query Scope Inconsistency
Appendix B Target Checking Module
Figure 7: Target Checking Module
Appendix C Training Settings
Parameter Value
per_device_train_batch_size 1
gradient_accumulation_steps 8
learning_rate 1.0e-4
num_train_epochs 2.0
lr_scheduler_type cosine
lora_rank 16
Table 4: Training hyperparameter configurations.
Appendix D Figures
Figure 8: SQL Probes in Error Cause Identification Stage
Appendix E Prompt used by SDE-SQL in the training-free setting
For approaches that do not rely on supervised training, it becomes particularly crucial to carefully craft and design prompt templates that can effectively guide the model to carry out Self-Driven Exploration behaviors in a controlled and meaningful manner. In this section, we present a comprehensive set of prompt templates that are utilized across different stages of the SDE-SQL pipeline to support this capability. Due to limitations in available space, certain detailed elements and specific prompt examples have been omitted, but the essential structures and core ideas are fully retained.
Figure 9: Prompt Template of Candidates Exploration
Figure 10: Prompt Template of Combinations Exploration
Figure 11: Prompt Template of Zero-shot Generation
Figure 12: Prompt Template of Solution Exploration
Figure 13: Prompt Template of Final Refinement
Figure 14: Prompt Template of Target Checking